OpenStudy (anonymous):

Which expression is equivalent to (1 – sin2 x)(sec x)?

OpenStudy (anonymous):

\[\sin^{2}\theta+\cos^{2}\theta=1\] \[1-\sin^{2}\theta=\cos^{2}\theta\] \[\sec \theta=\frac{ 1 }{ \cos \theta }\] \[(\cos^{2}\theta)(\frac{ 1 }{ \cos \theta })=\cos \theta\]

OpenStudy (anonymous):

thanks